# Complete Security and Privacy for AI Inference in Decentralized Systems

### v1

@@ 2.1.9 Zero-Knowledge Proofs for Foundation Models
2.1.10 zkDPS: Zero-Knowledge Decentralized Proof System at Nesa
At Nesa, we build a zero-knowledge system, zkDPS, for decentralized large language models. zkDPS is built upon zkLLM [13] and zkDL [11]. zkLLM [13] and zkDL [11] are state-of-the-art CUDA implementation of zero-knowledge proofs of machine learning. However, speed remains the most significant pain point for zero-knowledge proofs of foundation models. For example, it takes zkLLM roughly 15 minutes to generate a proof for the inference procedure of LLaMA-2 13B with an input token length of 2,048 [13]. In comparison, vanilla inference of zkLLM takes less than 0.1 seconds per token. Despite significant efforts have been made to accelerate generic zero-knowledge proof frameworks such as zk-STARK and zk-SNARK [23], they are still not fast enough, even much slower than zkLLM for billion-scale models. Therefore, the acceleration of zero-knowledge proofs is urgent for real-world applications.
@@
Another idea is to leverage the hardware acceleration of modern GPUs. zkLLMs has already used CUDA implementation to parallelize the computations in the proof generation. For example, the CUDA kernels such as Pedersen commitment, elliptic curve, and lookup tables have been re-implemented or adapted in zkLLMs [13, 11], making it the fastest implementation of zero-knowledge large language models at the submission of this paper. In Nesa, we are optimizing those kernels by engineering methods such as Tensor Parallelism and KV cache.
2.2 Sequential Vector Encryption
At Nesa, we have designed an innovative encryption method called Sequential Vector Encryption (SVE) which is a robust scheme inspired by our diagrammatic interpretation of homomorphic encryption applied on select linear layers rather than the full model. In particular, we focus on the modern LLM architecture but this framework can be applied to other types of neural networks such as CNNs. We consider the LLM to be a sequence of vector space operators. As mentioned previously, this sequence contains both linear and non-linear operators. In principle, the (intermediate) output vector of every operator contains information about both the input and the output of the model. We sometimes refer to this operator output as a (latent) representation vector, as it is a representation of the input data after (partial) processing through some intermediate computational operations in the neural network. We wish to apply SVE when there is some extra concern that an attacker can access these vectors and, subsequently, can extract sensitive information attributed to either the initial model input or model output. Our method focuses on obfuscating the output of every operator in the sequence of the LLMs so that any observation of the neural representation of the data in any intermediate latent space is not usable for this purpose. This extra level of security is particularly relevant for the use of models that heavily rely on open-source foundation models for which the intermediate latent spaces are available for extensive study by any actor.
The general technique for SVE is based on the simple idea of randomly transforming the outputs of every operator so that the intermediate vector representations can no longer be interpretable using a given method developed for the original model’s vector representations. In order to be usable for the original model, these representations need to be transformed into the original representation space before feeding to the next operator. Equivalently, we can transform the operator itself to the operator on the transformed vector space in the correct way, instead of transforming the vector itself, which is how the technique is implemented in practice as this allows efficient secret sharing - to be discussed below. In either case, the computational complexity of the transformation is the same. As the intermediate representations have dimensionalities on the order of roughly $10^{3}$ components (for reference a small LM such as BERT typically uses 768 dimensions and Llama 2 uses 4096 dimensions), any attempt to recover the transformed vectors means determining the parameters of the transformation (representable as a matrix), which are on the order of $10^{6}$ floating point parameters for a typical transformation. This would amount to solving a very large undetermined system.
There are two main considerations to put this into practice: (1) an appropriate transformation of each operator, and (2) a secret sharing method so that transformation matrices are never revealed during operation. The former (1) ensures that the transformations preserve the computational outputs of the original model and the latter (2) is required so that no agent can recover the encrypted representation vectors. The following will explain how we address both considerations.
Appropriately transforming each operator depends on the type of operator under consideration, for a linear operation, it can be transformed as a product of (invertible) transformation matrices and the original matrix representation of that operator (recoverable from the original LLM). For non-linear operators, we can take advantage of certain symmetries of the operator. For instance, the ReLU $R$ operator is pointwise, and thus $R$ commutes with any vector space permutation $P$ according to the following diagram:
[…]
Here we are transforming between real vector spaces of the same dimension. Since $P$ is always invertible, this diagram tells us that we can represent the ReLU operator as $R^{\prime}=P^{-1}RP$. This transformation alone is not necessarily enough to obscure the transformation as a bad actor with knowledge of the layer could in principle run a known vector through the open source model and compare the two representation vectors component-wise to figure out the transformation. Therefore we can scale with a random vector with all positive entries, as the $R$ also fits into the following commutative diagram:
[…]
where $M_{a}$ is multiplication by $a\in\mathbb{R}^{n}_{>0}$, a diagonal matrix with positive diagonal terms. Concatenating the two transformations we obtain the representation $R^{\prime\prime}=M_{a}^{-1}R^{\prime}M_{a}=M_{a}^{-1}P^{-1}RPM_{a}$, noting that any diagonal matrix with non-zero diagonal entries is invertible. As an aside, it is interesting to note that the number of per- mutations on an n-dimensional vector are simply n!, so even for BERT, the number of possible permutations is 768! whereas it is thought that there are roughly 60! atoms in the universe. Other nonlinear operators can be transformed using similar, but more careful, considerations. We finally package our transformation $T=PM_{A}$ as a single matrix $T$, with corresponding inverse $T=M_{a}^{-1}P^{-1}$. However, revealing this matrix directly in the procedural code distributed to agents will compromise security. Thus we use a form of secret sharing to obfuscate the transformation matrices.
To see how secret sharing can be incorporated in SVE, it is illuminating to consider how to transform a linear/nonlinear pair of operators as this is the basic structure of most neural network layers, including LLMs. We represent the single (original) $i$th layer $B_{i}$ as
${B_{i}=RA},$ (27)
where $A$ is a linear transformation and $R$ is our nonlinear ReLU as above. Consider $\overline{A}=SA$, where $S$ is the appropriate transformation matrix, and $\overline{R}=T^{-1}RT$, where $T$ is defined above. Then
$\begin{split}\overline{B_{i}}&=\overline{R}\overline{A}\\ &=(T^{-1}RT)(SA)\\ &=T^{-1}(RTSA)\\ &=T^{-1}RC,\end{split}$ (28)
where $C=TSA$ is a new matrix transformed from the original matrix $A$. Here we assume the output of the previous block is transformed according to $S^{1}$, and the output of $\overline{B_{i}}$ will be equivalent to the output of $B_{i}$ transformed by $T^{-1}$. Now the transformation pair $S$ and $S^{-1}$ cannot be recovered by the above agent but the transformation $T^{-1}$ is recoverable, which allows $S$ to also be recoverable with (public) knowledge of $A$. Thus by looking ahead to the next layer $B_{i+1}$ we can insert the appropriate transformation pair $U$ and $U^{-1}$ by absorbing $U^{-1}$ in to the linear representation for $B_{i+1}$ and using the representation
$\overline{B_{i}}=UT^{-1}RC=GRC,$ (29)
where $G=UT^{-1}$. Now the layer $B_{i}$ as $GRC$ appropriately transforms the layer so that none of the matrices $T^{-1}$, $U$, $T$ or $S$ can be recovered.
A final attack can be made if two agents in sequence collude. We mitigate this by injecting additional transformations in the sequential chain of operators, similar to how the transformation $G$ was constructed. These additional operators are sent, along with the encrypted representation vector at this point in the sequence to run on either a trusted third party such as a TEE, or on the original client to sequentially encrypt the representation vector. Thus, the encrypted outputs are always passed back to the holder of this additional transformation. This topology is natural in the sense of a decentralized network this can be performed at the same time as consensus is verified.
3 Security and Privacy for General Inference
Figure 4: Security flowchart for general inference, where Nesa designs CBV3.1) for model verification and leverages SL3.2) for data encryption and user privacy protection.
@@ 3 Security and Privacy for General Inference
In general inference scenarios, the results of AI inference tasks are less critical, allowing for faster processing speeds and lower costs without compromising basic security and privacy standards. These tasks include everyday activities such as checking the weather, recommending products, or filtering spam emails. Our strategy for ensuring model verification and integrity in these scenarios is Consensus-based Verification Check (CBV), which leverages the collective agreement of multiple nodes to ensure the correctness and integrity of model execution without revealing sensitive data. This approach is detailed in Section 3.1. For data privacy, we employ split learning (SL) [14, 26], a method where the model is divided into segments, and each segment is trained on different nodes to maintain data privacy by ensuring no single node has access to the complete dataset. More information on this technique can be found in Section 3.2. Fig. 4 provides an overview of the security flow of general inference at Nesa which first applies SL to encrypt the first (few) layers to “encrypt” the raw data, where the AI models are sharded and deployed to nodes with redundancy. For the nodes who have the same shard, their outputs will be verified by consensus for integrity.
3.1 Consensus-based Verification Check for Model Integrity
@@
Given the computational and scalability challenges associated with ZKML for verifying the integrity of LLMs in decentralized systems, Nesa proposes a consensus-based distribution verification (CDV) strategy for general inference scenarios. This strategy leverages the collective agreement of multiple nodes to ensure the correctness and integrity of model execution without revealing sensitive data.
@@
Taking Model Sharding into Account. In Nesa’s decentralized system, where ML models’ computational graphs may be sharded across multiple nodes for scalability, each node $i$ possesses a unique shard $\mathcal{M}_{i}$ of the complete model $\mathcal{M}$. This partitioning requires a specialized approach to Consensus-based Verification to accommodate the fragmented nature of model execution.
@@
    Model Reconstruction. After each shard has been independently verified, the shard-specific consensus results $\{y_{\text{con},1},y_{\text{con},2},\ldots,y_{\text{con},k}\}$ are combined to reconstruct the final model output $Y_{\text{final}}$. This comprehensive output can ensure the integrity of the complete model execution.
@@
Through its implementation, CDV offers a powerful solution to the challenges of verifying the integrity of distributed ML models in Nesa’s decentralized framework. By focusing on distributional characteristics rather than discrete output values, CDV not only elevates the verification process but also aligns with the goals of enhancing model security and maintaining stringent privacy standards.
3.2 Data Privacy Protection via Split Learning
Recognizing the challenges posed by encrypting data for use in decentralized inference systems, Nesa adopts Split Learning (SL) as a pragmatic solution to facilitate secure and efficient computation on encrypted data [28, 29]. Traditional encryption methods such as HE, while securing data at rest and in transit, render it costly for direct computation by obscuring its format and structure. This limitation is particularly problematic for processing with LLMs within a decentralized framework, where data privacy cannot be compromised.
@@ 3.2 Data Privacy Protection via Split Learning
It is noted that by using differential privacy with SL, the privacy will be improved at the cost of inference quality [32]. Thus, in Nesa’s framework, this is defined as a tunable parameter to be decided, given the user requirements.
By leveraging Split Learning, Nesa effectively navigates the complexities of data encryption within its decentralized inference system for LLMs. This approach not only preserves the confidentiality and integrity of user data but also ensures the operational feasibility of complex model computations, demonstrating a sophisticated balance between privacy preservation and computational pragmatism.
4 NTEE: Nesa’s Trusted Execution Environments for Security and Privacy
Trusted Execution Environments (TEEs) effectively establish security and privacy within Nesa’s decentralized inference architecture. Using TEEs, we can create secure isolation zones within the network’s nodes, guaranteeing the privacy and accuracy of data and computational operations. In Nesa’s scenarios, two key pieces of information should be protected from the node runners: (i) users’ input data for AI inference and (ii) private model parameters, where TEE provides a hardware-based solution for it.
@@ 4 NTEE: Nesa’s Trusted Execution Environments for Security and Privacy
By leveraging TEEs, Nesa can provide robust security and privacy measures that are essential for both critical and general inference scenarios, ensuring that sensitive data and model parameters are always protected from unauthorized access and tampering. Thus, we design a specialized version of TEEs, named NTEE, that best suits Nesa’s scenarios with decentralized inference. Note that NTEE is feasible for all node runners with or without GPUs.
@@
  • Low Overhead and High Performance: TEEs provide a secure computing environment with minimal performance overhead (i.e., can be as small as 3% as reported in different literature [33, 34]), ensuring efficient execution of AI inference tasks. This allows Nesa to maintain high performance in both critical and general inference scenarios.
@@
  • Elimination of Algorithm-Based Approaches: Utilizing TEEs negates the need for algorithm-based security measures like ZKML and HE discussed above in §2.1 and 2.2. This results in faster performance due to the inherent efficiency of hardware-based security, making it ideal for Nesa’s decentralized AI inference tasks.
  • Scalable and Secure Collaboration: TEEs enable secure collaboration across multiple nodes by creating isolated enclaves that protect data during computation. This is particularly beneficial for Nesa’s decentralized AI inference, allowing multiple organizations to pool their data securely for training and inference while maintaining data privacy and compliance.
4.1 Implementation of NTEE in Details
Nesa, as a decentralized AI provider, aims to have multiple node runners collaboratively leverage their hardware to perform AI inference for users. In this setup, Nesa employs one orchestrator within the Rich Execution Environment (REE) to facilitate the entire process within NTEE. The REE, unlike the TEE, is the standard, general-purpose operating environment that handles regular applications and processes without the enhanced security measures of a TEE. We have the user data $\mathbf{X}$ and the AI model $M$ (e.g., Llama 2 [30]), and we have $k$ node runners for hosting TEEs. Fig. 5 provides a high-level overview of our implementations of NTEE.
@@ 4.1 Implementation of NTEE in Details
Figure 5: Overview of Nesa’s NTEE paradigm for decentralized AI inference.
Step 1: TEE Initialization. Nesa first sends a blank Virtual Machine (VM) TEE to each node runner. This VM does not contain any data or model information initially, ensuring no sensitive information is at risk during the setup phase.
Step 2: Remote Attestation. The orchestrator initiates ”remote attestation” with NVIDIA to verify that the node runner’s machine meets the requirements for NVIDIA’s TEE. Remote attestation involves the following steps:
@@
Step 3: Secure Data and Model Transmission. Once attestation confirms the TEE’s security, Nesa securely transmits the user data $\mathbf{X}$ and the AI model fragment $M_{i}$ to the node runner’s TEE via an encrypted channel, similar to SSH:
@@
By implementing TEEs in this manner, Nesa ensures that the AI inference process is secure, protecting both the user’s data and the AI model throughout the entire lifecycle of the computation. This approach leverages the robustness of hardware-based security measures, providing high performance and strong guarantees of confidentiality and integrity. TEEs also facilitate mutual attestation, where both the user and the node runner can verify each other’s integrity, further enhancing the trustworthiness of the decentralized AI system.
4.2 Nesa’s Innovation of NTEE
Nesa operates in a unique decentralized AI setting, distinct from the general usage of TEEs where only one REE interacts with one TEE to run a model or a specific task. Nesa’s system involves multiple TEEs across various nodes to run a shard of an AI model, requiring advanced strategies to ensure efficiency and security.
Innovation 1: Communication Optimization Among Multiple TEEs. Nesa optimizes communication by minimizing data transfer between the orchestrator (considered as a REE) and each TEE. Instead of routing all communications through the orchestrator, we establish pre-set secure communication channels among the TEEs. This allows TEEs to communicate directly with each other, significantly reducing latency and improving overall system efficiency. Direct communication between TEEs is secured using encrypted channels to ensure data privacy and integrity during transit, and the owners of the nodes cannot see the communication either. This approach reduces bottlenecks and enhances the speed of decentralized inference processes, which is crucial for real-time applications.

### v2

@@ A.1.9 Zero-Knowledge Proofs for Foundation Models
A.1.10 zkDPS: Zero-Knowledge Decentralized Proof System
We build a zero-knowledge system, zkDPS, for decentralized large language models. zkDPS is built upon zkLLM [12] and zkDL [10]. zkLLM [12] and zkDL [10] are state-of-the-art CUDA implementation of zero-knowledge proofs of machine learning. However, speed remains the most significant pain point for zero-knowledge proofs of foundation models. For example, it takes zkLLM roughly 15 minutes to generate a proof for the inference procedure of LLaMA-2 13B with an input token length of 2,048 [12]. In comparison, vanilla inference of zkLLM takes less than 0.1 seconds per token. Despite significant efforts have been made to accelerate generic zero-knowledge proof frameworks such as zk-STARK and zk-SNARK [23], they are still not fast enough, even much slower than zkLLM for billion-scale models. Therefore, the acceleration of zero-knowledge proofs is urgent for real-world applications.
@@
Another idea is to leverage the hardware acceleration of modern GPUs. zkLLMs has already used CUDA implementation to parallelize the computations in the proof generation. For example, the CUDA kernels such as Pedersen commitment, elliptic curve, and lookup tables have been re-implemented or adapted in zkLLMs [12, 10], making it the fastest implementation of zero-knowledge large language models at the submission of this paper. We are optimizing those kernels by engineering methods such as Tensor Parallelism and KV cache.
Appendix B Security and Privacy for General Inference
@@
In general inference scenarios, the results of AI inference tasks are less critical, allowing for faster processing speeds and lower costs without compromising basic security and privacy standards. These tasks include everyday activities such as checking the weather, recommending products, or filtering spam emails. Our strategy for ensuring model verification and integrity in these scenarios is Consensus-based Verification Check (CBV), which leverages the collective agreement of multiple nodes to ensure the correctness and integrity of model execution without revealing sensitive data. This approach is detailed in Section B.1. For data privacy, we employ split learning (SL) [13, 15], a method where the model is divided into segments, and each segment is trained on different nodes to maintain data privacy by ensuring no single node has access to the complete dataset. More information on this technique can be found in Section B.2. Fig. 2 provides an overview of the security flow of general inference, which first applies SL to encrypt the first (few) layers to “encrypt” the raw data, where the AI models are sharded and deployed to nodes with redundancy. For the nodes who have the same shard, their outputs will be verified by consensus for integrity.
B.1 Consensus-based Verification Check for Model Integrity
@@
Given the computational and scalability challenges associated with ZKML for verifying the integrity of LLMs in decentralized systems, We proposes a consensus-based distribution verification (CDV) strategy for general inference scenarios. This strategy leverages the collective agreement of multiple nodes to ensure the correctness and integrity of model execution without revealing sensitive data.
@@
Taking Model Sharding into Account. In our decentralized system, where ML models’ computational graphs may be sharded across multiple nodes for scalability, each node $i$ possesses a unique shard $\mathcal{M}_{i}$ of the complete model $\mathcal{M}$. This partitioning requires a specialized approach to Consensus-based Verification to accommodate the fragmented nature of model execution.
@@
    Model Reconstruction. After each shard has been independently verified, the shard-specific consensus results
    $\{y_{\text{con},1},y_{\text{con},2},\ldots,y_{\text{con},k}\}$ are combined to reconstruct the final model output $Y_{\text{final}}$. This comprehensive output can ensure the integrity of the complete model execution.
@@
Through its implementation, CDV offers a powerful solution to the challenges of verifying the integrity of distributed ML models in our decentralized framework. By focusing on distributional characteristics rather than discrete output values, CDV not only elevates the verification process but also aligns with the goals of enhancing model security and maintaining stringent privacy standards.
B.2 Data Privacy Protection via Split Learning
Recognizing the challenges posed by encrypting data for use in decentralized inference systems, we adopts Split Learning (SL) as a pragmatic solution to facilitate secure and efficient computation on encrypted data [27, 28]. Traditional encryption methods such as HE, while securing data at rest and in transit, render it costly for direct computation by obscuring its format and structure. This limitation is particularly problematic for processing with LLMs within a decentralized framework, where data privacy cannot be compromised.
@@ B.2 Data Privacy Protection via Split Learning
It is noted that by using differential privacy with SL, the privacy will be improved at the cost of inference quality [31]. Thus, in our framework, this is defined as a tunable parameter to be decided, given the user requirements.
By leveraging Split Learning, we effectively navigates the complexities of data encryption within its decentralized inference system for LLMs. This approach not only preserves the confidentiality and integrity of user data but also ensures the operational feasibility of complex model computations, demonstrating a sophisticated balance between privacy preservation and computational pragmatism.
Appendix C Our Specialized Trusted Execution Environments for Security and Privacy
Trusted Execution Environments (TEEs) effectively establish security and privacy within our decentralized inference architecture. Using TEEs, we can create secure isolation zones within the network’s nodes, guaranteeing the privacy and accuracy of data and computational operations. In our scenarios, two key pieces of information should be protected from the node runners: (i) users’ input data for AI inference and (ii) private model parameters, where TEE provides a hardware-based solution for it.
@@ Appendix C Our Specialized Trusted Execution Environments for Security and Privacy
By leveraging TEEs, we can provide robust security and privacy measures that are essential for both critical and general inference scenarios, ensuring that sensitive data and model parameters are always protected from unauthorized access and tampering. Thus, we design a specialized version of TEEs that best suits our scenarios with decentralized inference. Note that our specialized TEE is feasible for all node runners with or without GPUs.
@@
  • Low Overhead and High Performance: TEEs provide a secure computing environment with minimal performance overhead (i.e., can be as small as 3% as reported in different literature [32, 33]), ensuring efficient execution of AI inference tasks. This allows us to maintain high performance in both critical and general inference scenarios.
@@
  • Elimination of Algorithm-Based Approaches: Utilizing TEEs negates the need for algorithm-based security measures like ZKML and HE discussed above in §A.1. This results in faster performance due to the inherent efficiency of hardware-based security, making it ideal for our decentralized AI inference tasks.
  • Scalable and Secure Collaboration: TEEs enable secure collaboration across multiple nodes by creating isolated enclaves that protect data during computation. This is particularly beneficial for our decentralized AI inference, allowing multiple organizations to pool their data securely for training and inference while maintaining data privacy and compliance.
@@
Figure 3: Overview of our specialized TEE paradigm for decentralized AI inference.
C.1 Implementation of Our Specialized TEE in Detail
As a decentralized AI provider, we aim to have multiple node runners collaboratively leverage their hardware to perform AI inference for users. In this setup, we employ one orchestrator within the Rich Execution Environment (REE) to facilitate the entire process within our specialized TEE. The REE, unlike the TEE, is the standard, general-purpose operating environment that handles regular applications and processes without the enhanced security measures of a TEE. We have the user data $\mathbf{X}$ and the AI model $M$ (e.g., Llama 2 [29]), and we have $k$ node runners for hosting TEEs. Fig. 3 provides a high-level overview of the implementations of our specialized TEE.
Step 1: TEE Initialization. We first send a blank Virtual Machine (VM) TEE to each node runner. This VM does not contain any data or model information initially, ensuring no sensitive information is at risk during the setup phase.
Step 2: Remote Attestation. The orchestrator initiates "remote attestation" with NVIDIA to verify that the node runner’s machine meets the requirements for NVIDIA’s TEE. Remote attestation involves the following steps:
@@ C.1 Implementation of Our Specialized TEE in Detail
Step 3: Secure Data and Model Transmission. Once attestation confirms the TEE’s security, we securely transmit the user data $\mathbf{X}$ and the AI model fragment $M_{i}$ to the node runner’s TEE via an encrypted channel, similar to SSH:
@@
By implementing TEEs in this manner, we ensure that the AI inference process is secure, protecting both the user’s data and the AI model throughout the entire lifecycle of the computation. This approach leverages the robustness of hardware-based security measures, providing high performance and strong guarantees of confidentiality and integrity. TEEs also facilitate mutual attestation, where both the user and the node runner can verify each other’s integrity, further enhancing the trustworthiness of the decentralized AI system.
C.2 Our Innovations in Specialized TEE
Our system operates in a unique decentralized AI setting, distinct from the general usage of TEEs where only one REE interacts with one TEE to run a model or a specific task. Our system involves multiple TEEs across various nodes to run a shard of an AI model, requiring advanced strategies to ensure efficiency and security.
Innovation 1: Communication Optimization Among Multiple TEEs. We optimize communication by minimizing data transfer between the orchestrator (considered as a REE) and each TEE. Instead of routing all communications through the orchestrator, we establish pre-set secure communication channels among the TEEs. This allows TEEs to communicate directly with each other, significantly reducing latency and improving overall system efficiency. Direct communication between TEEs is secured using encrypted channels to ensure data privacy and integrity during transit, and the owners of the nodes cannot see the communication either. This approach reduces bottlenecks and enhances the speed of decentralized inference processes, which is crucial for real-time applications.